The sum of the two digits of a two digit number is 15 and the difference between the two digits of the two digit number is 3. What is the product of the two digits of the two-digit number?
Answer by lwsshak3(11628)

The sum of the two digits of a two digit number is 15 and the difference between the two digits of the two digit number is 3. What is the product of the two digits of the two-digit number?
***
let u=units digit
let t=tens digit
..
u+t=15
u-t= 3
add
2u=18
u=9
t=15-u=6
product of the two digits of the two-digit number=9*6=54